OpenStudy (anonymous):

Determine the radius of a circle whose arc length measures 57.1 meters and central angle measures 1.46 radians. a. 39 b. 102 c. 83 d. 59

OpenStudy (jasmineflvs):

hello

OpenStudy (jasmineflvs):

can you help me?

OpenStudy (anonymous):

hi and sure I'll try if you return the favor as well

OpenStudy (jasmineflvs):

im not that high in math im asking a question for my friend

OpenStudy (jasmineflvs):

the graph y > x^2-6x +7 contains which points

OpenStudy (anonymous):

is x^2 means radical 2? i wanna get clear on that

OpenStudy (jasmineflvs):

i think it means squared

OpenStudy (jasmineflvs):

plzz help

OpenStudy (anonymous):

okay so substitute y=0 so the equation should be 0=x^2-6x+7

OpenStudy (jasmineflvs):

okay anything else

OpenStudy (anonymous):

i hope this is what you're looking for

OpenStudy (jasmineflvs):

k thanks

OpenStudy (anonymous):

mhm good luck!

OpenStudy (anonymous):

so the arclenght is equal to 2pi*radius, im not sure how the radians come into play

OpenStudy (jasmineflvs):

okay just a sec it was supposed to be the graph y= x^2-6x +7 contains which points and the ^ after the x means squared

Directrix (directrix):

@Jasmineflvs What are the points which are options to the question: y= x^2-6x +7

OpenStudy (anonymous):

do you know the website wolframalpha? i used that a lot for my math hw

OpenStudy (jasmineflvs):

idk Directrix

OpenStudy (jasmineflvs):

here they are A. (0, 0) B. (-3, 2) C. (5, -1) D. (3, -1)

OpenStudy (jasmineflvs):

can you answer that for me it is for a friend

OpenStudy (anonymous):

wait is this a different question?

OpenStudy (jasmineflvs):

yes

OpenStudy (anonymous):

hmmm not sure, don't you just plug in the points on that equation? I have to go sorry.

OpenStudy (jasmineflvs):

Directrix do you have the answer

OpenStudy (anonymous):

but you shoyld check out the site wolframalpha.com

OpenStudy (jasmineflvs):

okay he needs an answer though

Directrix (directrix):

I suggest we test points. I am working on what to post to help you see the answer. Just a sec.

OpenStudy (jasmineflvs):

okay

OpenStudy (jasmineflvs):

cause he needs it soon

Directrix (directrix):

y= x^2-6x +7 Let's test the point (3,-1) Substitue 3 for x and -1 for y. IS -1 > 9 - 6*3 + 7 -1 > -9 + 7 Yes or No @Jasmineflvs

OpenStudy (jasmineflvs):

idk i am just doing this for a friend ill ask

OpenStudy (jasmineflvs):

thank you

Directrix (directrix):

You are welcome.